OpenStudy (kellyspeakslouder):

Find the inverse of the function. f(x) = 4x3 - 7 A. Not a one-to-one function B. f-1(x) = cube root of x minus divided by four plus seven C. f-1(x) = cube root of quantity x plus seven divided by four D. f-1(x) = cube root of quantity x minus seven divided by four

OpenStudy (kellyspeakslouder):

@triciaal

OpenStudy (kellyspeakslouder):

@triciaal

OpenStudy (twwc960):

To find the inverse of f(x)=4x^3-7, first write y=4x^3-7 then interchange x and y and solve for y. So for the inverse, x=4y^3-7, 4y^3=x+7 or y^3=(x+7)/4, which means y= the cube root of (x+7)/4, so the answer is C.

OpenStudy (kellyspeakslouder):

@triciaal

OpenStudy (kellyspeakslouder):

@zepdrix

OpenStudy (kellyspeakslouder):

pleaseeee help meeee

OpenStudy (triciaal):

follow the directions from earlier switch the x and y then make y the subject of the equation remember to keep the balance

OpenStudy (triciaal):